ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 196, 2013

Adamaszek, M. Clique complexes and graph powers  295

Almeida, J. and A. Costa Presentations of Schützenberger groups of minimal subshifts

1

Andrews, U. and T. Makuluni The index set of uncountably categorical theories

491

Barthelmé, T. A natural Finsler–Laplace operator  375
Benanti, F. On the exponential growth of graded Capelli polynomials  51

Bergen, J., D. Riley and H. Usefi Lie superalgebras whose enveloping algebras satisfy a non-matrix polynomial identity

161

Braun, B., J. Browder and S. Klee Cellular resolutions of ideals defined by nondegenerate simplicial homomorphisms

321

Browder, J. See B. Braun Costa, A. See J. Almeida Dzhafarov, D. D. and C. Mummert On the strength of the finite

intersection principle

345

Gallardo-Gutiérrez, E. A. and J. R. Partington Norms of composition operators on weighted Hardy spaces

273

Ganchev, G. and V. Milousheva Timelike surfaces with zero mean curvature in Minkowski 4-space

413

Ginzburg, V. L. and D. Hein The Arnold conjecture for Clifford symplectic pencils

95

González-Avilés, C. D. Abelian class groups of reductive group schemes

175

Gurel-Gurevich, O. and R. Peled Poisson thickening  215

Hein, D. See V. L. Ginzburg Jorgensen, P. E. T. and E. P. J. Pearse A discrete Gauss–Green

identity for unbounded Laplace operators, and the transience of random walks

113

Klee, S. See B. Braun Komjáth, P. A note on chromatic number and connectivity of infinite

graphs

499

Komjáth, P. The list-chromatic number of infinite graphs  67

Makuluni, T. See U. Andrews Matsumoto, K. A certain synchronizing property of subshifts and

flow equivalence

235

Milousheva, V. See G. Ganchev Mummert, C. See D. D. Dzhafarov

Nitzan, N. and M. A. Perles Radon stability  483

Partington, J. R. See E. A. Gallardo-Gutiérrez Pearse, E. P. J. See P. E. T. Jorgensen Peled, R. See O. Gurel-Gurevich Perles, M. A. See N. Nitzan Riley, D. See J. Bergen Roy, R. and P. Schmid Steinitz classes for Galois extensions of

Dedekind rings

285

Salehi Golsefidy, A. Lattices of minimum covolume are nonuniform

363

Schmid, P. See R. Roy

Speight, G. Surfaces meeting porous sets in positive measure  435

Sznajdman, J. A residue calculus approach to the uniform Artin–Rees lemma

33

Usefi, H. See J. Bergen

Wang, Yu Lie superhomomorphisms on Lie ideals in superalgebras  461

Ye, S. Erratum to “A unified approach to the plus-construction, Bousfield localization, Moore spaces and zero-in-the-spectrum examples”

507